OpenStudy (anonymous):

solve by the substitution method 6x+7y=-27 and x=7-5y

OpenStudy (saifoo.khan):

x = -8 y = 3

OpenStudy (anonymous):

what is the solution of the system

OpenStudy (anonymous):

suntitutimg eq 2 in eq 1 ===> 6(7-5y)+7y=-27 ===> 42-30y+7y = -27 ===> y = 69/23 = 3

OpenStudy (anonymous):

===> x=7-15 = -8